number theory a number is divided by 5, then 3 is added to the quotient. after subtracting 10, the result is 30. what is the number.
Answer by mananth(16946)

let the number be x
(x/5 +3)-10=30
(x/5+3)=40
x/5=40-3
x/5=37
x=5*37
x=185